OpenStudy (anonymous):

Find the quadratic polynomial ax^2+bx+c which best fits the function f(x)=8^x at x=0, in the sense that g(0)=f(0), and f'(0)=g'(0), and f''(0)=g''(0). g(x)=?

OpenStudy (anonymous):

First of all calculate the f(x), f'(x) and f''(x) values for the given function \[f(x)=8^x\]\[f \prime(x)=8^x\ln8\]\[f \prime \prime(x)=8^x\ln^28\]Now we can find values for x=0\[f(0)=1\]\[f \prime(0)=\ln8\]\[f \prime \prime(0)=\ln^28\]Now consider the polynomial g(x) and it's derivatives\[g(x)=ax^2+bx+c\]\[g \prime(x)=2ax+b\]\[g \prime \prime(x)=2a\]We should also find these when x is 0\[g(0)=c\]\[g \prime(0)=b\]\[g \prime \prime(0)=2a\]As stated that f''(0)=g''(0) we can replace g''(0) with f''(0). So we get\[f \prime \prime(0)=2a\]\[\ln^28=2a\]\[a=(\ln^28)/2\]Repeat this process to determine the values for b and c\[b=\ln8\]\[c=1\]Now simply substitute these values into the original polynomial g(x) and you will get your desired function.

OpenStudy (anonymous):

Thanks for the problem, really made me think. Anyone feel free to correct me if I made a mistake. :)

OpenStudy (anonymous):

thank you for your help! it is correct i just checked